Macroeconomics economics

Suppose that a market is described by the following supply and demand equations:

QS = 2P

QD = 300 – P

a.       Solve for the equilibrium price and the equilibrium quantity.


b.      Suppose that a tax of T is placed on buyers, so the new demand equation is

QD = 300 – (P + T).

Solve for the new equilibrium. What happens to the price received by sellers, the price paid by buyers, and the quantity sold?

 

c.       Tax revenue is T X Q. Use your answer to part (b) to solve for tax revenue as a function of T. Graph this relationship for T between 0 and 300.


d.      The deadweight loss of a tax is the area of the triangle between the supply and demand curves. Recalling that the area of a triangle is 1 ⁄2 x base x height, solve for deadweight loss as a function of T. Graph this relationship for T between 0 and 300. (Hint: Looking sideways, the base of the deadweight loss triangle is T, and the height is the difference between the quantity sold with the tax and the quantity sold without the tax.)


e.       The government now levies a tax on this good of $200 per unit. Is this a good policy? Why or why not? Can you propose a better policy?
Answer #1

Answer a: Equilibrium quantity:

d = Q s

300-P= 2P

300= 3P

P= 100

Q= 2P= 2*100= 200units

Answer b: Now new Quantity demanded= 300-(P+T)

s = 2P

300-(P+T)= 2P

300-T= 3P

P= 100- T/3

New equilibrium price that seller received= 100-T/3

Equilibrium price that buyers paid = P+T = 100-T/3+T

= 100+2T/3

Quantity sold = 2P=2(100-T/3)= 200-2T/3


answered by: brownbarbara